40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Tue Feb 22 07:53:06 2022, Yehonathan, Update, WOPO, Waking up WOPO 
    Reply  Fri Feb 25 17:09:53 2022, shruti, Update, WOPO, Waking up WOPO - green beam 
       Reply  Wed Mar 2 14:38:55 2022, Yehonathan, Update, WOPO, Waking up WOPO - some more fiddling and a plan 
          Reply  Fri Mar 4 11:49:45 2022, shruti, Update, WOPO, Waking up WOPO - attempts at readout LO_shot_noise.pdf
             Reply  Thu Mar 10 17:12:01 2022, awade, Update, WOPO, Waking up WOPO - attempts at readout 
                Reply  Wed Mar 16 12:22:44 2022, Yehonathan, Update, WOPO, Waking up WOPO - attempts at readout Moku_Noise.pdfPD_Connected_no_light.pdfDiff_Channel.pdf
                   Reply  Tue Mar 22 09:19:38 2022, shruti, Update, WOPO, Waking up WOPO - back to green path before_opt.jpgaround_waveguide.png
Message ID: 2723     Entry time: Tue Feb 22 07:53:06 2022     Reply to this: 2725
Author: Yehonathan 
Type: Update 
Category: WOPO 
Subject: Waking up WOPO 

{Shruti, Yehonathan}

On Friday, we came down to QIL to poke around the WOPO setup. The first thing we noticed is that the setup on the wiki page is obsolete and in reality, the 532nm light is coming directly from the Diablo module.

There were no laser goggles for 532nm so we turned on the 1064nm (Mephisto) only. The pump diode current was ramped to 1A. We put a power meter in front of Mephisto and opened the shutter. Rotating the HWP we got 39mW. We dialed it back so that 5mW is coming out of the polarizer.

The beam block was removed. We disconnected the LO fiber end from the fiber BS - there is light coming out! we connected a power meter to the fiber end using an FC/PC Fiber Adapter Plate. The power read 0.7mW. By aligning the beam into the LO fiber we got up to 3.3mW.

We connected the BHD PDs to the scope on the table to observe the subtraction signal. Channel 2 was negative so we looked at the sum channel.

Time ran out. We ramped down the diode current and turned off Mephisto.

Next time we should figure out the dark current of the BHD and work toward observing the shot noise of the LO.

ELOG V3.1.3-